Suppose that X ~ fx (x10), where θ E Θ. Suppose that T = T(X) is a sufficient statistic for θ. Prove the following statement: If W-W(X) is the uniformly minimum variance ụnbi. ased estimator (UMVUE) of 0, then W-E(WT) with probability one.
Answer #1

The proof of the above theorem is given below.

んひalso an UE of withfiniteariame . Now, we note thot , the conditonal dist WIT isindependent ob Hence, E (HIT) i iadependext o e& in a unchon of T. otthatis, E(WIT) is a statistic bases on T- Note that, θ = E(w) = EE(w/t) す E(HIT) an UE of 9, based on T, Again, as w「‘ UMWE var(h)Var (E(r) an otter UE ofe-. var(u)= var (E(UIT)) der witんprobability